A plane traveled 2992 miles with the wind in 5.5 hours and 2563 miles against the wind in the same amount of time. Find the speed of the plane in still air and the speed of the wind.
The speed of the plane in still air is
___ mph. ​(Simplify your​ answer.)

Answer by ankor@dixie-net.com(22740)

A plane traveled 2992 miles with the wind in 5.5 hours and 2563 miles against the wind in the same amount of time.
Find the speed of the plane in still air and the speed of the wind.
:
let s = the speed in still air
let w = the speed of the wind
then
(s+w) = effective speed with the wind
and
(s-w) = effective speed against the wind
:
Write a distance equation for each way. Dist = time * speed
5.5(s+w) = 2992
5.5(s-w) = 2563
Distribute
5.5s + 5.5w = 2992
5.5s - 5.5w = 2563
-------------------adding eliminates w, find s
11s + 0 = 5555
s = 5555/11
s = 505 mph is the speed of the plane in still air